OpenStudy (anonymous):

The sum to n terms of the series \[1^2-2^2+3^2-4^2+5^2-6^2........\]

OpenStudy (anonymous):

My approach: (1+(n-1)2)^2 = (2n-1)^2 (2+(n-1)2)^2 = (2n)^2 4n^2-4n+1 - 4n^2 = -4n+1 \[\sum_{1}^{n}-4k+1 \]

OpenStudy (anonymous):

=- 2 (n(n+1) + n -2n^2 - 2n + n = -2n^2 - n

OpenStudy (experimentx):

is it finite series or infinite series

OpenStudy (experimentx):

Hint: + and - terms differently and try it.

OpenStudy (anonymous):